Nueva Company reported the following pretax data for its first year of operations.

Net sales 7,700
Cost of goods available for sale 5,970
Operating expenses 1,818
Effective tax rate 17 %
Ending inventories:
If LIFO is elected 726
If FIFO is elected 888


How much more will Nueva report in income tax if it elects FIFO instead of LIFO?

Multiple Choice

  • $28.

  • $120.

  • $109.

  • $92.
Answer #1

Calculate how much more will Nueva report in income tax

Difference in ending inventories = 726-888 = 162

Much more = 162*17% = 28

So answer is a) $28